the perimeter of a rectangle is 152m. The length is 4m more than twice the width. find the demensions
Found 2 solutions by checkley77, isharkk:
Answer by checkley77(12844)

152=2L+2W
L=2W+4
152=2(2W+4)+2W
152=4W+8+2W
152=6W+8
6W=152-8
6W=144
W=144/6
W=24 ANS. FOR THE WIDTH.
L=2*24+4
L=48+4
L=522ANS. FOR THE LENGTH.
PROOF:
152=2*52+2*24
152=104+48
152=152

Answer by isharkk(3)

152=2L+2W
L=2W+4
152=2(2W+4)+2W
152=4W+8+2W
152=6W+8
6W=152-8
6W=144
W=144/6
W=24 ANS. FOR THE WIDTH.
L=2*24+4
L=48+4
L=522ANS. FOR THE LENGTH.
PROOF:
152=2*52+2*24
152=104+48
152=152 D